OpenStudy (anonymous):

Solve S = 3πr + 7q for r. Answer r = the quantity S minus seven q, divided by three pi r = the quantity S plus seven q, divided by three p r = S + 7q – 3p r = S – 3p – 7q

OpenStudy (anonymous):

S = 3πr + 7q for r. 3nr = s - 7q -> r = s/3n - 7q/ 3n

OpenStudy (anonymous):

hmmm you sure

OpenStudy (anonymous):

yea yea your right

OpenStudy (anonymous):

r = the quantity S minus seven q, divided by three pi

OpenStudy (anonymous):

ok